OpenStudy (anonymous):

1/5 (7y-3) + 1/20 (12y+21)

OpenStudy (anonymous):

1/5(7y-3)+1/20(12y+21 7y/5-3/5+12y/20+21/20 28y/20+12y/20-12/20+21/20 40y/20-9/20 2y-9/20 2y=9/20 y=9/40

OpenStudy (anonymous):

wait so after i get 40y+9 /20 how do i reduce it